# Graphon Mixtures

### v1

@@ 6 Estimating $U$ for infinite number of partitions
Figure 2: Unique log degree values of a $(U,W)$ mixture graph $G_{n_{i}}$. The line fitted to points $\{(j,\log(\text{deg }_{G_{n_{i}}}v_{(j)})\}_{j=1}^{k_{i}}$ is shown in red and the line fitted to degrees generated from $W$ is shown in green. The line fitted to points $\{(j,\log\frac{m_{s_{i}}}{j})\}_{j=1}^{N}$ is shown in blue.
When $U$ has infinite partitions, Proposition 5.3 does not hold. In the finite partition case, there is a widening gap between the hub degrees generated by $U$ and the degrees generated by $W$. But in the infinite partition case, there are always smaller values $p_{i}$ that fill this gap. However, we show that the degrees generated by $U$ are different to those generated by $W$. Let $k_{i}$ denote the number of hub nodes in $G_{n_{i}}$ generated by $U$ with degrees higher than those generated by $W$. Our goal is to estimate $k_{i}$. Thus, we focus on the larger degrees of $G_{n_{i}}$ and consider the unique degree values of $G_{n_{i}}$ greater than some percentile $C$. We consider the ordered, unique log degree values $(j,\log(\text{deg }_{G_{n_{i}}}v_{(j)})$, where $\text{deg }_{G_{n_{i}}}v_{(j)}$ is the $j$th highest degree in $G_{n_{i}}$. To these log degree values we fit two lines using the OLS estimator. We show that the line fitted to the first $k_{i}$ points $\{(j,\log(\text{deg }_{G_{n_{i}}}v_{(j)})\}_{j=1}^{k_{i}}$ has a steeper slope compared to the line fitted to the rest of the points, which are generated by $W$. In Figure 2 these two lines are shown in red and green respectively. We do this by comparing with the line fitted to points on the curve $y=\log(m_{s_{i}}/x)$, which is shown in blue in Figure 2.
@@ 7 Experiments
We conduct experiments with synthetic and real data on a standard laptop. As a baseline comparison method for degree prediction, we use scale-free graph properties discussed in Bollobás et al. (2001). Another simple comparison method detailed in Appendix F is used for estimating $U$. As the evaluation metric we use the Mean Absolute Percentage Error (MAPE) given by $100\times\frac{1}{N}\sum_{i=1}^{N}\frac{|\hat{y}_{i}-y_{i}|}{y_{i}}$. For degree prediction $y_{i}$ denotes the top-$k$ degree values and for estimating $U$, $y_{i}$ denotes the mass-partition elements $p_{i}$. More details of the experiments are given in Appendix F.
@@ 7.1 Illustration with synthetic data
With synthetic data experiments we focus on 3 tasks: (1) predict the top-$k$ degrees, (2) estimate $U$ when the mass-partition has finite non-zero elements and (3) estimate $U$ when the mass-partition has infinite non-zero elements. To predict the top-$k$ degrees we use training graphs with $n_{i}$ nodes and test graphs with $n_{j}$ nodes. Using equation (1) we predict the top-$k$ degrees for graphs in the test set. To estimate $U$ we only use training graphs. When $U$ has finite partitions we use equation (4) and for infinite partitions we use equation (6). For each task we conduct 4 experiments. For degree prediction we consider $W_{1}=\exp(-(x+y))$, $W_{2}=0.1$ and $U_{1}$ with mass-partitions $\bm{p}_{1}\propto\{1/j^{1.2}\}_{j=2}^{50}$ and $U_{2}$ with mass-partition $\bm{p}_{2}\propto\{1/1.2^{j}\}_{j=2}^{50}$. The mass-partitions are normalized so that they add up to 1. Table 1 gives the results of the experiments.
7.2 Citation graphs and social networks
For observed real world graphs, we do not have a ground truth model for $U$, hence we are only able to compute MAPE for predicted degrees. Using Hep-PH Physics citations dataset (Leskovec et al., 2005; Gehrke et al., 2003) and Facebook (FB) links dataset (Viswanath et al., 2009) we predict degrees. In Hep-PH dataset, each paper is a node and if paper $i$ cites paper $j$, then there is an edge between nodes $i$ and $j$. We focus on monthly graphs, where the graph is updated with new papers and new citations. We consider graphs for months 80, 85, 90, 95 and 100. Using each graph at month $\ell$, we predict the top 10 degrees of the graph at months $\ell+6$ up to $\ell+10$. Similarly, for the FB dataset we consider graphs on days 520, 540, 560, 580 and 600 with prediction steps 60 to 100, stepping by 10. Table 2 gives the results averaged over the prediction step and the graph index.
Table 2: Our proposed method is better than baseline for HEP-PH and FB datasets. Average MAPE over different prediction steps and graphs reported with standard deviation in parenthesis.
Dataset Prediction Step MAPE Graph Index MAPE
Proposed Baseline Proposed Baseline
Hep-PH 6 2.18 (1.68) 3.67 (2.25) 80 2.89 (1.93) 5.07 (3.12)
7 2.66 (2.03) 5.47 (2.58) 85 3.37 (2.95) 4.64 (2.60)
8 3.15 (2.32) 6.07 (2.98) 90 2.15 (1.72) 5.44 (2.80)
9 3.42 (2.62) 6.84 (3.16) 95 2.47 (1.59) 5.60 (3.21)
10 3.53 (2.53) 7.81 (3.33) 100 4.20 (2.63) 7.52 (4.01)
FB 60 7.13 (4.48) 10.5 (5.76) 520 9.12 (4.50) 13.1 (6.37)
70 7.84 (4.75) 12.5 (5.58) 540 12.5 (4.31) 17.8 (6.16)
80 8.09 (4.73) 13.1 (6.18) 560 7.82 (3.37) 13.9 (4.64)
90 8.56 (5.14) 15.3 (5.72) 580 4.73 (3.00) 10.7 (4.51)
100 8.06 (5.14) 15.8 (5.70) 600 4.53 (3.23) 10.7 (4.94)
@@ 8 Conclusions
We have modeled the highest degrees of sparse $(U,W)$ mixtures and estimated $U$ as this is the source of sparsity. We can estimate $U$ with high accuracy when $U$ has finite disjoint cliques. When $U$ has infinitely many disjoint cliques, the accuracy depends on the rate at which the clique proportions go to zero. We have focused on sparse $(U,W)$ mixtures. Dense $(U,W)$ mixtures supplement graphs generated by a single graphon $W$. While we briefly touch on this in Appendix B, this topic can be explored further. This work focused on estimating the sparse component $U$, which is the new piece in the puzzle. Many methods have been developed to estimate the classical graphon. Once the effect of $U$ is removed from the mixture, $W$ can be estimated from the remaining graph. This is an avenue for future work.
Supplement to Graphon Mixtures
@@ Appendix A Background: Limits of sparse graphs and line graphs
$W_{G}(x,y)=\begin{cases}1&\,\text{if}\quad ij\in E(G)\\ 0&\,\text{otherwise}\,,\end{cases}$
@@
[…]
where the supremum is taken over all measurable sets $A$ and $B$ of $[0,1]$.
@@ A.1 Square-degree property and max-degree condition
deg $\sum\text{deg }v_{i,n}^{2}\geq c_{1}\left(\sum\text{deg }v_{i,n}\right)^{2}\,.$
@@
deg $\frac{\sum\text{deg }v_{i,n}^{2}}{m^{2}}\geq\frac{d_{\max,n}^{2}}{m^{2}}\geq c^{2}>0\,,$ (8)
@@
$\liminf_{n\to\infty}\frac{\sum\text{deg }v_{i,n}^{2}}{m^{2}}=\lim_{n\to\infty}\frac{\sum\text{deg }v_{i,n}^{2}}{m^{2}}$ $=c>0$ (9)
@@
Noting the series deg $\frac{\sum\text{deg }v_{i,n}^{2}}{m^{2}}$ is absolutely convergent and the individual terms are positive $\left(\frac{d_{(i)}^{2}}{m^{2}}>0\right)$ we have
@@ Appendix B Mixtures of sparse and dense graphs
We provide an example motivating graphon mixtures. Consider the following example: Let $\{G_{n}\}_{n}$ be a sequence of graphs such that originally $G_{n}\sim G(n,p)$ where $G(n,p)$ denotes a graph generated from the Erdős–Rényi model where $G_{n}$ has $n$ nodes and the edge probability is $p$. Then we modify each graph in the sequence such that a randomly selected node in $G_{n}$ is connected to all other nodes in $G_{n}$. Figure 3 shows the empirical graphons for $n\in\{50,100,200,400\}$ with $p=0.1$ for this construction.
@@
This is an example of a dense ($U,W)$-mixture sequence, which converges to $W=0.1$ in the cut-metric (Definition A.4). While the mixture is dense, it has a sparse component that a single graphon $W$ does not fully explain. That is, graphs generated from $W=0.1$, would not have the persistent anomalous node shown in Figure 4. This shows that $W=0.1$ does not fully capture the observed graphs.
For this example, we can identify the sparse component using equation (3). While we give another example of dense mixtures in Section B.1 we do not delve into dense mixtures in this work. Our focus is on sparse $(U,W)$ mixtures.
B.1 A $(U,W)$ mixture example

### v2

@@ 6 Estimating $U$ for infinite number of partitions
Figure 3: Unique log degree values of a $(U,W)$ mixture graph $G_{n_{i}}$. The line fitted to points $\{(j,\log(\text{deg }_{G_{n_{i}}}v_{(j)}))\}_{j=1}^{k_{i}}$ is shown in red and the line fitted to degrees generated from $W$ is shown in green. The line fitted to points $\{(j,\log\frac{m_{s_{i}}}{j})\}_{j=1}^{N}$ is shown in blue.
When $U$ has infinite partitions, Proposition 5.3 does not hold. In the finite partition case, there is a widening gap between the hub degrees generated by $U$ and the degrees generated by $W$. But in the infinite partition case, there are always smaller values $p_{i}$ that fill this gap. However, we show that the degrees generated by $U$ are different to those generated by $W$. Let $k_{i}$ denote the number of hub nodes in $G_{n_{i}}$ generated by $U$ with degrees higher than those generated by $W$. Our goal is to estimate $k_{i}$. Thus, we focus on the larger degrees of $G_{n_{i}}$ and consider the unique degree values of $G_{n_{i}}$ greater than some percentile $C$. We consider the ordered, unique log degree values $\log(\text{deg }_{G_{n_{i}}}v_{(j)})$, where $\text{deg }_{G_{n_{i}}}v_{(j)}$ is the $j$th highest degree in $G_{n_{i}}$. To the points $(j,\log(\text{deg }_{G_{n_{i}}}v_{(j)}))$, we fit two lines using the OLS estimator. We show that the line fitted to the first $k_{i}$ points $\{(j,\log(\text{deg }_{G_{n_{i}}}v_{(j)}))\}_{j=1}^{k_{i}}$ has a steeper slope compared to the line fitted to the rest of the points, which are generated by $W$. In Figure 3 these two lines are shown in red and green respectively. We do this by comparing with the line fitted to points on the curve $y=\log(m_{s_{i}}/x)$, which is shown in blue in Figure 3.
@@ 7 Experiments
Table 2: Comparison of proposed method with Bollobás et al. (2001), Caron & Fox (2017) and 3 versions of Kronecker graphs (Leskovec et al., 2010) on 6 datasets. Average MAPE reported with standard deviation in parenthesis. Best results in bold. Sig. denotes statistical significance with $\alpha=0.1$.
Dataset Bollobas Caron-Fox Kronfit1 Kronfit2 Kronfit3 UW-Mixture Sig.
FB 0.130 (0.041) 0.175 (0.103) 0.289 (0.068) 111.0 (14.9) 0.518 (0.047) 0.101 (0.044)
HEP-PH 0.177 (0.105) 0.435 (0.066) 0.351 (0.165) 20.20 (4.01) 0.294 (0.115) 0.061 (0.046)
MOOC 0.015 (0.012) 0.854 (0.014) 0.772 (0.068) 0.797 (0.022) 0.835 (0.031) 0.005 (0.005)
SMS 0.051 (0.033) 0.426 (0.051) 0.702 (0.032) 0.747 (0.024) 0.029 (0.020)
UCI 0.201 (0.106) 0.757 (0.044) 0.404 (0.098) 2.19 (0.409) 0.908 (0.011) 0.185 (0.090)
Yahoo 0.030 (0.037) 0.446 (0.046) 0.356 (0.080) 0.834 (0.014) 0.021 (0.011)
We conduct experiments with synthetic and real data on a standard laptop. As the evaluation metric we use the Mean Absolute Percentage Error (MAPE) given by $100\times\frac{1}{N}\sum_{i=1}^{N}\frac{|\hat{y}_{i}-y_{i}|}{y_{i}}$. For degree prediction, $y_{i}$ denotes the top-$k$ degree values and for estimating $U$, $y_{i}$ denotes the mass-partition elements $p_{i}$. More details of the experiments are given in Appendix F.
@@ 7.1 Illustration with synthetic data
With synthetic data experiments we focus on 3 tasks: (1) predict the top-$k$ degrees, (2) estimate $U$ when $\bm{p}$ has finite partitions and (3) estimate $U$ when $\bm{p}$ has infinite partitions. To predict the top-$k$ degrees we use training graphs with $n_{i}$ nodes and test graphs with $n_{j}$ nodes. Using equation (1) we predict the top-$k$ degrees for graphs in the test set. To estimate $U$ we only use training graphs. When $U$ has finite partitions we use equation (4) and for infinite partitions we use equation (6). For each task we conduct 4 experiments. For degree prediction we consider $W_{1}=\exp(-(x+y))$, $W_{2}=0.1$ and $U_{1}$ with mass-partition $\bm{p}_{1}\propto\{1/j^{1.2}\}_{j=2}^{50}$ and $U_{2}$ with mass-partition $\bm{p}_{2}\propto\{1/1.2^{j}\}_{j=2}^{50}$. The mass-partitions are normalized so that they add up to 1. A baseline comparison is conducted using scale-free graph properties (Bollobás et al., 2001). Table 1 gives the results of the experiments.
7.2 Real-world networks
For observed real world graphs, we do not have a ground truth model for $U$, hence we only compute MAPE for predicted degrees. We use Facebook (FB) links (Viswanath et al., 2009), Hep-PH Physics citations (Leskovec et al., 2005; Gehrke et al., 2003), MOOC interactions (Kumar et al., 2019), SMS (Wu et al., 2010), UCI Messages (Panzarasa et al., 2009) and Yahoo messages (Rossi & Ahmed, 2015) datasets. Each dataset is given as an edgelist with timestamped edges. We use daily graphs for Facebook links, MOOC interactions, SMS, and UCI messages datasets. For Yahoo messages we use 2-hourly graphs and for Hep-PH citations monthly graphs. Each dataset consists of a sequence of growing graphs $\{G_{n_{i}}\}_{i}$ and we consider $G_{n_{i}}$ for $i\in\{20,\ldots,24\}$ as training graphs. We select the test graphs $G_{n_{j}}$ such that $j=\min_{k}\{|G_{n_{k}}|:|G_{n_{k}}|-|G_{n_{i}}|\geq 500\}$, i.e., $G_{n_{j}}$ is the first graph that has more than 500 nodes compared to $G_{n_{i}}$. We predict the top-10 degrees of $G_{n_{j}}$ using Bollobás et al. (2001), Caron & Fox (2017) and 3 variations of Kronecker graphs (Leskovec et al., 2010). Furthermore, we test for statistical significance between the top 2 methods. Table 2 gives the results with blanks denoting timed out instances. The $(U,W)$-mixture estimate is significantly better for 5 out of 6 datasets.
@@ 8 Conclusions
Our focus has been on $U$, the new piece in the puzzle. We have modeled the highest degrees of sparse $(U,W)$ mixtures and estimated $U$. We can estimate $U$ with high accuracy when $U$ has finite disjoint cliques. When $U$ has infinitely many disjoint cliques, the accuracy depends on the rate at which the clique proportions go to zero. We have focused on sparse $(U,W)$ mixtures. While we briefly touch on dense $(U,W)$ mixtures in Appendix B, this topic can be explored further. Exploring sparsity resulting from other types of graphs such as paths and rings is an avenue for future work.
@@ Appendix A Background: Limits of sparse graphs and line graphs
[…]
@@
$\lVert W\rVert_{\square}=\sup_{A,B}\left|\int_{A\times B}W(x,y)\,dxdy\right|\,$, (7)
where the supremum is taken over all measurable sets AA and $B$ of $[0,1]$.
@@ A.1 Square-degree property and max-degree condition
$\sum\text{deg }v_{i,n}^{2}\geq c_{1}\left(\sum\text{deg }v_{i,n}\right)^{2}\,.$
@@
$\frac{\sum\text{deg }v_{i,n}^{2}}{m^{2}}\geq\frac{d_{\max,n}^{2}}{m^{2}}\geq c^{2}>0\,$, (8)
@@
[…]
@@
Noting the series $\frac{\sum\text{deg }v_{i,n}^{2}}{m^{2}}$ is absolutely convergent and the individual terms are positive $\left(\frac{d_{(i)}^{2}}{m^{2}}>0\right)$ we have
@@ Appendix B Mixtures of sparse and dense graphs
B.1 $(U,W)$ mixture examples
We give two examples of $(U,W)$-mixture graphs in this section.
B.1.1 Example 1
We consider $U$ and $W$ shown in Figure 4 where $U$ is given by the mass-partition $\bm{p}=(0.5,0.3,0.2)$ and $W$ is given by a stochastic block model graphon. Graphon $W$ has 4 equally sized communities where the edge probability of two nodes $i$ and $j$ within the same community is given by $p_{ij}=0.6$ and the edge probability of two nodes $i$ and $k$ in two different communities is given by $p_{ik}=0.025$.
Figure 4: Graphons $U$ and $W$ used in Example 1.
Figures 6 and 6 show examples of mixture graphs from this $U$ and $W$. In Figure 6 the dense part $G_{d_{i}}$ has 100 nodes and the sparse part $G_{s_{i}}$ has 600 nodes. As $G_{s_{i}}$ has a larger number of nodes, the resulting mixture graph $G_{n_{i}}$ has large hubs, resembling a graph from a sparse sequence. In contrast, in Figure 6 the dense part $G_{d_{i}}$ has 100 nodes and the sparse part $G_{s_{i}}$ has 50 nodes making the mixture graph $G_{n_{i}}$ resemble a graph from a dense sequence.
From this $U$ and $W$ we construct a sparse $(U,W)$-mixture sequence with $n_{s_{i}}=\lceil n_{d_{i}}\times\sqrt{5i}\rceil$ for $i\in\{1,\ldots 20\}$. For each mixture graph $G_{n_{i}}$, we compute the global clustering coefficient, the mean distance of the graph and its spectral radius. Table 3 gives the results. We see that the clustering coefficient does not go to zero in this example.
Table 3: The global clustering coefficient, mean distance and spectral radius of the sparse $(U,W)$ mixture graphs sequence discussed in Example 1
$n_{d_{i}}$ $n_{s_{i}}$ Clustering Coefficient Mean Distance Spectral Radius
100 224 0.32 4.06 17.08
200 633 0.31 3.83 33.60
300 1162 0.33 3.67 50.50
400 1789 0.33 3.75 67.46
500 2500 0.33 3.66 84.18
600 3287 0.33 3.57 101.17
700 4142 0.33 3.54 117.99
800 5060 0.33 3.42 134.99
900 6038 0.33 3.48 151.80
1000 7072 0.33 3.48 168.85
1100 8158 0.33 3.47 186.21
1200 9296 0.33 3.41 203.01
1300 10481 0.33 3.46 219.61
1400 11714 0.33 3.37 236.22
1500 12991 0.33 3.32 252.36
1600 14311 0.33 3.28 270.54
1700 15674 0.33 3.27 287.00
1800 17077 0.33 3.25 303.76
1900 18519 0.33 3.29 320.70
2000 20000 0.33 3.26 337.55
Figure 5: A $(U,W)$-mixture graph with $n_{d_{i}}=100$ and $n_{s_{i}}=600$ where $U$ and $W$ are shown in Figure 4. The mixture graph $G_{n_{i}}$ resembles a graph from a sparse sequence.
Figure 6: A $(U,W)$-mixture graph with $n_{d_{i}}=100$ and $n_{s_{i}}=50$ where $U$ and $W$ are shown in Figure 4. The mixture graph $G_{n_{i}}$ resembles a graph from a dense sequence.
B.1.2 Example 2
Next we consider $U$ having the mass-partition $\bm{p}\propto\left(\frac{1}{2},\frac{1}{3},\ldots,\frac{1}{10}\right)$ such that $\sum_{i}p_{i}=1$. Let $W$ be given by
$W=0.9\exp\left(\frac{(-y^{2}-(x-1)^{2})}{0.05^{2}}\right)+0.9\exp\left(\frac{(-(y-1)^{2}-x^{2})}{0.05^{2}}\right)+0.9\exp\left(-\frac{\left(\sin\left(3\pi/4\right)x+\cos\left(3\pi/4\right)x\right)}{0.05^{2}}\right)\,$,
as in Xia et al. (2023). Graphon $W$ produces dense ring-type graphs. The two graphons are shown in Figure 7. Figure 9 shows the dense part $G_{d_{i}}$ sampled from $W$ with $n_{d_{i}}=100$, the sparse part $G_{s_{i}}$ with $n_{s_{i}}=100$ and the mixture graph $G_{n_{i}}$. Figure 9 shows the dense part $G_{d_{i}}$ with $n_{d_{i}}=100$, the sparse part $G_{s_{i}}$ with $n_{s_{i}}=600$ and the mixture graph $G_{n_{i}}$. We see that the hub structure is stronger in Figure 9 as $n_{s_{i}}>n_{d_{i}}$.
Again we construct a sparse mixture sequence from this $U$ and $W$. Similar to Example 1, we let $n_{s_{i}}=\lceil n_{d_{i}}\times\sqrt{5i}\rceil$ for $i\in\{1,\ldots 20\}$. For each mixture graph $G_{n_{i}}$, we compute the global clustering coefficient, the mean distance of the graph and its spectral radius. Table 4 gives the results. Similar to Example 1, we see a robust clustering coefficient as $n_{s_{i}}$ increases.
Table 4: The global clustering coefficient, mean distance and spectral radius of the sparse $(U,W)$ mixture graphs sequence discussed in Example 2
$n_{d_{i}}$ $n_{s_{i}}$ Clustering Coefficient Mean Distance Spectral Radius
100 224 0.27 6.06 10.62
200 633 0.29 4.51 21.84
300 1162 0.31 4.28 33.22
400 1789 0.31 4.23 44.12
500 2500 0.32 4.06 55.80
600 3287 0.32 4.14 66.76
700 4142 0.31 4.08 77.61
800 5060 0.32 4.09 89.42
900 6038 0.32 3.98 100.76
1000 7072 0.32 4.03 111.96
1100 8158 0.32 4.04 123.18
1200 9296 0.32 4.03 133.78
1300 10481 0.32 4.02 145.67
1400 11714 0.32 3.88 156.22
1500 12991 0.32 3.97 167.71
1600 14311 0.32 3.96 178.63
1700 15674 0.32 3.96 190.49
1800 17077 0.32 3.92 201.24
1900 18519 0.32 3.86 212.29
2000 20000 0.32 3.91 223.61
Figure 7: Graphons $U$ and $W$ used in Example 2.
Figure 8: A $(U,W)$-mixture graph with $n_{d_{i}}=100$ and $n_{s_{i}}=100$ where $U$ and $W$ are shown in Figure 7.
Figure 9: A $(U,W)$-mixture graph with $n_{d_{i}}=100$ and $n_{s_{i}}=600$ where $U$ and $W$ are shown in Figure 7
We see that the mixture graphs $G_{n_{i}}$ shown in Figures 6 and 6 are quite different from those in Figures 9 and 9. The main reason is that graphon $W$ stipulates the dense part of the mixture and as such, different $W$ produces different mixture graphs.
B.2 An example where the cut metric cannot explain the anomalous node
@@
Consider the following example: Let $\{G_{n}\}_{n}$ be a sequence of graphs such that originally $G_{n}\sim G(n,p)$ where $G(n,p)$ denotes a graph generated from the Erdős–Rényi model where $G_{n}$ has $n$ nodes and the edge probability is $p$. Then we modify each graph in the sequence such that a randomly selected node in $G_{n}$ is connected to all other nodes in $G_{n}$. Figure 10 shows the empirical graphons for $n\in\{50,100,200,400\}$ with $p=0.1$ for this construction.
@@
The empirical graphons (Definition A.2) converge to $W=0.1$ in the cut-metric (Definition A.4). While the graph sequence is dense, it has a sparse component that a single graphon $W$ does not fully explain. That is, graphs generated from $W=0.1$, would not have the persistent anomalous node shown in Figure 11. This shows that $W=0.1$ does not fully capture the observed graphs. Modeling these graphs as a mixture gives the flexibility to account for the persisting high degree node.
B.3 Generating dense and sparse graph sequences with a $(U,W)$ mixture